# Classical and quantum chaos of closed strings on a charged confining holographic background

### v1

@@ 1 Introduction
Various works have analyzed chaos in QCD observables in different holographic settings. The study of chaos in QCD chiral condensates in $\mathcal{N}=2$ supersymmetric QCD with $SU(N_{c})$ gauge group at large $N_{c}$ and ’t Hooft coupling was done in  [23] and the $\mathcal{N}=2$ theory was found to be chaotic. Similarly, a chaos-based mapping of the holographic QCD phase diagram in linear sigma models was obtained in [24]. Using a hanging open string approach, the chaos analysis in the quark-antiquark pair in the QCD deconfined phase was carried out in [25]. Their work was extended to analyze the anisotropic effects of the magnetic field and chemical potential on chaotic properties of QCD in top-down and bottom-up holographic settings in [26, 27, 28, 29, 30]. Additionally, tools like out-of-time-ordered correlator (OTOC), Pole-skipping, and Krylov complexity have been suggested as order parameters for the deconfinement transition [31, 32, 33]. One must use a confining background geometry to study chaos in the dual confined phase. This is usually done by taking the AdS soliton background, having a cigar-shaped geometry, in the gravity side [34, 35]. The geometry proves instrumental in investigating confinement within holographic gauge theories, as it encapsulates key characteristics such as the mass gap and confinement, phenomena commonly observed in low-temperature QCD. Both open and closed strings have been utilized to analyze QCD chaos in the confined phase by making use of AdS soliton backgrounds [36, 37, 38, 39]. Open strings, whose endpoints represent quark-antiquark pairs, offer insight into quark confinement and meson dynamics, while closed strings, which correspond to glueballs in the dual QCD theory, allow for the exploration of the gluonic sector [40, 41, 42, 43]. For other related works on chaos in open and closed strings, see [44, 45, 46, 47, 48, 49, 50, 51, 52, 53] for a necessarily biased selection.
In a series of earlier works [54, 55], the dynamics of closed strings in an uncharged AdS soliton background were examined using classical chaos diagnostics, such as Poincaré sections and Lyapunov exponents, as well as quantum chaos measures, including the analysis of energy level-spacing distributions. It was found that the bosonic strings on an AdS soliton background are nonintegrable and exhibit chaos. The uncharged AdS soliton is simply a double analytic continuation of the Schwarzschild AdS black hole solution and can be thought of as the ground state of the $\mathcal{N}=4$ supersymmetric Yang-Mills theory. This is interesting considering that the AdS soliton corresponds to a confined phase in the dual field theory side, and the nonintegrable nature of the closed string dynamics in the AdS soliton background accordingly provides valuable information about the chaotic behavior of the glueball spectrum of the dual confined phase.
@@ 2 Charged AdS solution and closed string dynamics
$\mathcal{S}=\frac{1}{16\pi G}\int d^{5}x\sqrt{-g}\left(R+\frac{12}{\ell^{2}}-\frac{3}{4}\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}\right)+\frac{1}{16\pi G}\int d^{5}x\sqrt{-g}\mathcal{F}\wedge\mathcal{F}\wedge\mathcal{A}\,,$ (2.1)
@@
$\mathcal{L}=\frac{r^{2}}{\ell^{2}}\left(-\dot{t}^{2}+f(r)\dot{\phi}^{2}+\dot{x}^{2}-x^{2}\alpha^{2}\right)+\frac{\ell^{2}}{r^{2}f(r)}\dot{(}r)^{2},$ (2.8)
@@ 4.1 Level-spacing distribution
A reference plot for level-spacing distributions is given in [55] for a particle in a box with a deformed potential in the nonintegrable case. The distribution functions agree with Poisson and Wigner-Dyson level statistics for the integrable and chaotic cases, respectively. The peculiar structure of spacing between adjacent levels in non-integrable systems arises from the fact that the corresponding energies show quantum mechanical level repulsion, meaning they are spaced apart. Therefore, due to level repulsion, small energy differences are suppressed. In the integrable case, eigenvalues from distinct sectors are interdependent, leading to an absence of level repulsion.
@@ 4.5 Numerical analysis, discussion, and results
Figure 14: Early time microcanonical OTOCs for different values of $n$ and $k$. Here first 200 eigenvalues and eigenfunctions were taken during the calculations. The results are divided into two categories: low energy ($n=5,10,20$) and high energy ($n=110,120,150$). There is a visible effect of charge for low-energy microcanonical OTOCs in the view of lowering the early exponential growth rate, while for the high-energy case, we only get aperiodic oscillations and no exponential growth.
@@ 5 Conclusion
Our overall investigation here not only demonstrates the interplay of charge and energy in shaping the chaotic behavior of the closed string across both classical and quantum domains in the confined phase but also provides a new window for further studies. For instance, the transition from chaotic dynamics, marked by level repulsion at lower energies, to an integrable regime as energy increases underscores a complex energy dependence that merits further in-depth investigation of the intermediate energy regimes, which could potentially clarify the subtle transitions between chaos and integrability. However, the full string spectrum remains an open question, especially in the context of a complete quantum theory without the simplifying assumptions of minisuperspace quantization. While the minisuperspace technique offers useful insights, it inherently limits our understanding of the complete dynamical behavior of the system. In the absence of minisuperspace quantization, the challenge lies in determining the entire spectrum. This invites investigation into non-perturbative methods and new theoretical frameworks that can account for the full spectrum. Our work can also be extended to understand the onset of quantum chaos in other holographic QCD models, in particular, in bottom-up models which are known to be relatively more useful in modeling QCD holographically.

### v2

@@ 1 Introduction
Various works have analyzed chaos in QCD observables in different holographic settings. The study of chaos in QCD chiral condensates in $\mathcal{N}=2$ supersymmetric QCD with $SU(N_{c})$ gauge group at large $N_{c}$ and ’t Hooft coupling was done in  [23] and the $\mathcal{N}=2$ theory was found to be chaotic. Similarly, a chaos-based mapping of the holographic QCD phase diagram in linear sigma models was obtained in [24]. Using a hanging open string approach, the chaos analysis in the quark-antiquark pair in the QCD deconfined phase was carried out in [25]. Their work was extended to analyze the anisotropic effects of the magnetic field and chemical potential on chaotic properties of QCD in top-down and bottom-up holographic settings in [26, 27, 28, 29, 30]. Additionally, tools like out-of-time-ordered correlator (OTOC), Pole-skipping, and Krylov complexity have been suggested as order parameters for the deconfinement transition [31, 32, 33]. One must use a confining background geometry to study chaos in the dual confined phase. This is usually done by taking the AdS soliton background, having a cigar-shaped geometry, in the gravity side [34, 35]. The geometry proves instrumental in investigating confinement within holographic gauge theories, as it encapsulates key characteristics such as the mass gap and confinement, phenomena commonly observed in low-temperature QCD. Both open and closed strings have been utilized to analyze QCD chaos in the confined phase by making use of AdS soliton backgrounds [36, 37, 38, 39]. Open strings, whose endpoints represent quark-antiquark pairs, offer insight into quark confinement and meson dynamics, while closed strings, which correspond to glueballs in the dual QCD theory, allow for the exploration of the gluonic sector [40, 41, 42, 43]. For other related works on chaos in open and closed strings, see [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] for a necessarily biased selection.
In a series of earlier works [55, 56], the dynamics of closed strings in an uncharged AdS soliton background were examined using classical chaos diagnostics, such as Poincaré sections and Lyapunov exponents, as well as quantum chaos measures, including the analysis of energy level-spacing distributions. It was found that the bosonic strings on an AdS soliton background are nonintegrable and exhibit chaos. The uncharged AdS soliton is simply a double analytic continuation of the Schwarzschild AdS black hole solution and can be thought of as the ground state of the $\mathcal{N}=4$ supersymmetric Yang-Mills theory. This is interesting, considering that the AdS soliton corresponds to a confined phase in the dual field theory side, and the nonintegrable nature of the closed string dynamics in the AdS soliton background accordingly provides valuable information about the chaotic behavior of the glueball spectrum of the dual confined phase.
@@ 2 Charged AdS solution and closed string dynamics
$\mathcal{S}=\frac{1}{16\pi G}\int d^{5}x\sqrt{-g}\left(R+\frac{12}{\ell^{2}}-\frac{3}{4}\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}\right)+\frac{1}{16\pi G}\int\mathcal{F}\wedge\mathcal{F}\wedge\mathcal{A}\,,$ (2.1)
@@
$\mathcal{L}=\frac{r^{2}}{\ell^{2}}\left(-\dot{t}^{2}+f(r)\dot{\phi}^{2}+\dot{x}^{2}-x^{2}\alpha^{2}\right)+\frac{\ell^{2}}{r^{2}f(r)}\dot{r}^{2},$ (2.8)
@@ 4.1 Level-spacing distribution
A reference plot for level-spacing distributions is given in [56] for a particle in a box with a deformed potential in the nonintegrable case. The distribution functions agree with Poisson and Wigner-Dyson level statistics for the integrable and chaotic cases, respectively. The peculiar structure of spacing between adjacent levels in non-integrable systems arises from the fact that the corresponding energies show quantum mechanical level repulsion, meaning they are spaced apart. Therefore, due to level repulsion, small energy differences are suppressed. In the integrable case, eigenvalues from distinct sectors are independent, leading to an absence of level repulsion.
@@ 4.5 Numerical analysis, discussion, and results
Figure 14: Early time microcanonical OTOCs for different values of $n$ and $k$. Here first 200 eigenvalues and eigenfunctions were taken during the calculations. The results are divided into two categories: low energy ($n=5,10,20$) and high energy ($n=110,120,150$). The color coding is done as: red for $q=0.0$, green for $q=0.3$, cyan for $q=0.6$ and orange for $q=0.9$. There is a visible effect of charge for low-energy microcanonical OTOCs in the view of lowering the early exponential growth rate, while for the high-energy case, we only get aperiodic oscillations and no exponential growth.
@@ 5 Conclusion
Our overall investigation here not only demonstrates the interplay of charge and energy in shaping the chaotic behavior of the closed string across both classical and quantum domains in the confined phase but also provides a new window for further studies. For instance, the transition from chaotic dynamics, marked by level repulsion at lower energies, to an integrable regime as energy increases underscores a complex energy dependence that merits further in-depth investigation of the intermediate energy regimes, which could potentially clarify the subtle transitions between chaos and integrability.
However, the full string spectrum remains an open question, especially in the context of a complete quantum theory without the simplifying assumptions of minisuperspace quantization. The minisuperspace approximation provides a viable framework for probing quantum chaos in the closed string spectrum, despite inherent truncations of the full string degrees of freedom. This approach retains only the center-of-mass dynamics and a subset of spatial modes, neglecting fermionic fields, background fluxes, and oscillator sectors of the full theory, as noted in studies of other confining geometries [74]. While some level of systematic errors are expected due to this approximation, its validity in capturing universal spectral properties is supported by precedents across quantum cosmology, string theory, and lattice QCD. The minisuperspace methodology, initiated in quantum cosmology [91], simplifies dynamics by restricting to symmetric configurations, arguing these dominate the path integral. Similarly, in Liouville theory [92], truncating to minisuperspace preserves critical features like critical exponents despite omitting oscillator modes. For highly symmetric systems, such as AdS3 string configurations [93], this approximation reproduces exact spectral results in integrable sectors. Crucially, as shown in [94, 95], even truncated dynamics retain key statistical signatures (e.g., level repulsion) when the classical limit is chaotic, aligning with studies where minisuperspace results match lattice QCD observations of hadronic spectral statistics [16, 96]. This indicates that while absolute energies may deviate, the universality of level spacing distributions (GOE vs. Poisson) might remain robust. This resilience mirrors findings in Liouville theory and related studies, where high symmetry allows the approximation to capture dominant spectral features with reasonable accuracy [92, 94].
A rigorous quantification of deviations from the full quantum theory remains an open challenge. Future work could address this by comparing minisuperspace results with full-string spectra in solvable limits, such as near-integrable deformations [95]. Nonetheless, the ability of the minisuperspace framework to reproduce chaotic signatures, which is supported by string-theoretic precedents and lattice QCD parallels, underscores its utility as an important and relevant tool for studying quantum chaos in holographic confining gauge theories, despite its simplified nature.